# Tuning Colloidal Reactions

### v1

@@ Acknowledgements.
We thank Sam Schoenholz for his work developing JAX-MD and the members of the Brenner Group at Harvard University for helpful discussions. This material is based upon work supported by the Office of Naval Research (ONR N00014-17-1-3029, ONR N00014-23-1-2654), the NSF Grant DMR-1921619, the NSF AI Institute of Dynamic Systems (#2112085), and a grant from the Simons Foundation (#1141499).References
@@
- Murugan et al. [2015a] A. Murugan, J. Zou, and M. P. Brenner, Undesired usage and the robust self-assembly of heterogeneous structures, Nature communications 6, 6203 (2015a).
- Zeravcic et al. [2014] Z. Zeravcic, V. N. Manoharan, and M. P. Brenner, Size limits of self-assembled colloidal structures made using specific interactions, Proceedings of the National Academy of Sciences 111, 15918 (2014).
@@
- Keim et al. [2019] N. C. Keim, J. D. Paulsen, Z. Zeravcic, S. Sastry, and S. R. Nagel, Memory formation in matter, Reviews of Modern Physics 91, 035002 (2019).
- Murugan et al. [2015b] A. Murugan, Z. Zeravcic, M. P. Brenner, and S. Leibler, Multifarious assembly mixtures: Systems allowing retrieval of diverse stored structures, Proceedings of the National Academy of Sciences 112, 54 (2015b).
- Holmes-Cerfon et al. [2013] M. Holmes-Cerfon, S. J. Gortler, and M. P. Brenner, A geometrical approach to computing free-energy landscapes from short-ranged potentials, Proceedings of the National Academy of Sciences 110, E5 (2013).
- Holmes-Cerfon [2016] M. C. Holmes-Cerfon, Enumerating rigid sphere packings, siam REVIEW 58, 229 (2016).
- Murugan and Vaikuntanathan [2017] A. Murugan and S. Vaikuntanathan, Topologically protected modes in non-equilibrium stochastic systems, Nature communications 8, 13881 (2017).

### v2

@@ Acknowledgements.
We thank Sam Schoenholz for his work developing JAX-MD and the members of the Brenner Group for helpful discussions. This work is supported by the Office of Naval Research (ONR N00014-17-1-3029, ONR N00014-23-1-2654), the NSF Grant DMR-1921619, the NSF AI Institute of Dynamic Systems (#2112085), and a grant from the Simons Foundation (#1141499). E.M.K. acknowledges support from a Simons Foundation Junior Fellowship under Grant Number 1141499.References
@@
- Caspar and Klug [1962] D. L. Caspar and A. Klug, Physical principles in the construction of regular viruses, in Cold Spring Harbor symposia on quantitative biology, Vol. 27 (Cold Spring Harbor Laboratory Press, 1962) pp. 1–24.
- Twarock and Luque [2019] R. Twarock and A. Luque, Structural puzzles in virology solved with an overarching icosahedral design principle, Nature communications 10, 4414 (2019).
@@
- Matthies et al. [2023] M. Matthies, R. Krueger, A. Torda, and M. Ward, Differentiable partition function calculation for rna, bioRxiv , 2023 (2023).